OpenStudy (anonymous):

Find f(a), f(a+h), and the difference quotient f(a+h) - (fa)/h where h =/= 0.

OpenStudy (anonymous):

What is f?

OpenStudy (anonymous):

The function.

OpenStudy (anonymous):

So what is your question :-D.

OpenStudy (anonymous):

lol you asked didn't you?

OpenStudy (anonymous):

Oops forgot to put f(x) = 1/(x+1)

OpenStudy (anonymous):

:-D

OpenStudy (anonymous):

It is a derived

OpenStudy (anonymous):

do you have an "a" as well or is it just "a"

OpenStudy (anonymous):

It's just "a". I'm getting \[((1/a+h+1)-(1/a+1))/h \] for f(a+h) - (fa)/h where h =/= 0 but that simplifies to no solution. I'm not sure if that's correct.

OpenStudy (anonymous):

\[ \frac{f(x+h)-f(x)}{h} = \frac{\frac{1}{1+x+h}-\frac{1}{1+x}}{h} = \frac{1+x - (1+x+h)}{(1+x+h)(1+x)h} = -\frac{1}{(1+x+h)(1+x)} \]

OpenStudy (anonymous):

it is correct so far but you need some algebra

OpenStudy (anonymous):

ythemanifold cleared the fractions, then combined like terms in the numerator and canceled the h from top and bottom

OpenStudy (anonymous):

also note that you should have -h in the numerator, which explains the "-" sign out front at the last step